# FoodLMM: A Versatile Food Assistant using Large Multi-modal Model

### v1

@@ 5.3 FoodVQA Results
Our FoodVQA consists of three tasks, i.e., food classification, ingredient recognition, and recipe generation. For these tasks, we compare the performance of our model with baseline models that achieves state-of-the-art performances. Table 3 summarizes the results. As shown in the table, for all these three tasks, our FoodLMM outperforms the SOTA methods and achieves the best results. Specifically, for food classification, by fine-tuning on Food-101 dataset, our model attains an accuracy at 93.93%, which is 3.07% higher than the SOTA method [68]. For ingredient recognition, our model outperforms CACLNet [59], a SOTA method on VIREO Food-172 dataset, for 3.2% in terms of F1. For recipe generation, without any extra information, our FoodLMM significantly outperforms FIRE [17], the SOTA method on Recipe1M dataset that leverages ground-truth ingredient labels for recipe generation. Figure 3) further shows qualitative examples for foodVQA. As can be seen, our FoodLLM can generate accurate answers for food classification, ingredient recognition as well as recipe generation, showing the effectiveness of our model.
@@ 5.4 Nutrition Estimation Results
We further evaluate the performance of our model in nutrition estimation, the results are shown in Table 3. After training Stage 1, FoodLMM has been able to accurately estimate the overall nutritional value based on food images, decreasing the prediction error by $4.5\%$ on average compare to the previous SOTA. Example of using FoodLMM to estimate the total nutritional value of food is shown in Figure3. Moreover, FoodLMM is also able to predict the nutritional value for specific ingredient in dishes, which is named as Referring Nutrition Estimation. It is worthwhile mentioning that previous methods lack the ability of Referring Nutrition Estimation. For evaluating the performance of Referring Nutrition Estimation, we select the top three ingredients with the largest mass as referred ingredients. Table 3 summarizes the results. In the table, refer@$k$st denotes referring the ingredient with the $k$th highest mass. The results show that although further fine-tuning does not enhance the performance of overall nutrition estimation, it significantly improves the accuracy of nutrition estimation for specific ingredients.
@@ 5.5 Referring Segmentation Results
For referring segmentation, we consider three scenarios: One-to-One, One-to-Many and One-to-Zero [39]. One-to-one means that only one single ingredient being referred in each query, and the model should provide a single mask. One-to-many indicates that multiple ingredients are referred each time, and the model should generate corresponding distinct masks. One-to-zero means to refer objects not exist in the image, the model should not give any mask.
One-to-one. We compare our FoodLMM with LISA [48] on FoodSeg103 and UECFoodPIX datasets, the results are shown in Table 3. For standard referring segmentation (one-to-one), FoodLMM’s cIoU scores significantly exceed those of LISA, achieving 20% improvement (0.65 to 0.78) on FoodSeg103 and 25% (0.72 to 0.90) on UECFoodPIX datasets. Since we deliberately train FoodLMM’s one-to-zero ability, it is inevitable that the model may give a zero mask to the correct referring, i.e., mistakenly identifying the queried ingredient as non-existent. We exam the probability of accurate response (ACC) to measure this impact. Notably, the decrease in FoodLMM’s accuracy is minimal (99.16% on FoodSEG103 and 98.69% on UECFoodPIX) despite significant improvements in cIoU.
One-to-many. We then evaluate the performance of segmentation involving multiple ingredients referred in a query. Table  3 also summarizes the segmentation results of three ($refer@3$) and five ($refer@5$) referred ingredients. The results show that FoodLMM, after the first training stage (FoodLMM S1), achieves high cIoU scores (around 0.7) on the $refer@3$ task on two datasets, while also maintaining high accuracy. However, as the number of ingredients to be referred increased, the segmentation become more challenging. Nevertheless, the fine-tuned model (FoodLMM FT) attain higher cIoU scores in all scenarios and maintain a high probability of accurate responses (over 99%) even in complex segmentation tasks. Figure 3 shows the examples of referred ingredient segmentation.
One-to-zero To further examine the reasoning performance of our FoodLLM, we evaluate its performance under a challenging scenario, where the model is required to segment the ingredients that are not present in the image. For such query, the model should reject to return the segmentation mask. The performance of one-to-zero segmentation are shown in Table 3. As LISA is not trained with one-to-zero samples, the accuracy in identifying absent ingredients on both FoodSeg103 and UECFoodPIX datasets is 0.0%. In contrast, our FoodLMM, attain the accuracy of 73.07% on FoodSeg103 and 77.50% on UECFoodPIX, when handle one single nonexistent referring ($refer@1$), after the first train stage. The task becomes more challenging when there are three absent referring (refer@3) in the query. Nevertheless, FoodLMM still maintains a high accuracy, surpassing 80% after fine-tuning. Figure 3 presents an example where the user requests FoodLMM to segment watermelon from the seafood vegetable soup, but the model declines.
@@ 5.6 Performance on the proposed Benchmarks
Multi-turn Conversation. We further evaluate the performance of our model on Multi-turn Conversation dataset. The evaluation is performed by user study. Specifically, we invited 10 participants to manually evaluate the satisfaction score (1 to 5) of 20 sets of multi-turn dialogues (totalling 200 dialogue sets, all of the questions are randomly selected from the test set of our FoodDialogues). The results are summarized in Table 4. The majority of the dialogues (173 out of 200) receive high scores (>>>=3), and 4 is the most frequent score. It indicates that our FoodLMM is capable of closely aligning with user preferences and generating high-quality answers that are both accurate and explanatory.
@@
Reasoning Segmentation. We compare the performance of our LLM with LISA in reasoning segmentation on our FoodReasonSeg benchmark. The results are shown in Table 5. From the results, LISA achieves a quite low gIoU and cIoU, which are $0.19$ and $0.17$, respectively. It indicates that although LISA is excellent in the general domain’s reasoning segmentation, due to lacking specialized knowledge in the food domain, its performance in food reasoning segmentation is less impressive. In contrast, our FoodLMM S1 achieves much better performance by learning food knowledge from public datasets. After the Stage 2 fine-tuning, the performance of FoodLMM is significantly improved, with the gIoU score increasing from $0.35$ to $0.71$ and the cIoU score from $0.34$ to $0.72$ compared to FoodLMM S1, indicates that FoodLMM acquired more expertise and stronger reasoning ability from the constructed datasets.
@@ 7.1 Stage 1: Public Food Datasets
Figure 4: Templates for Food Classification task.
Food Classification. We adopt the VIREO Food-172 to construct the instruction-following data for the food classification task. Our 10 question templates are detailed in Figure 4.
Figure 5: Templates for Ingredient Recognition task.
Ingredient Recognition. We use VIREO Food-172 and Recipe1M to develop instruction-following data for the ingredient recognition task. The templates are outlined in Figure 5.
Figure 6: Templates for Recipe Generation task.
Recipe Generation. We construct instruction-following data for the recipe generation task using recipes from the Recipe1M dataset that have corresponding images. Moreover, consistent with [80], we map the 16k ingredients in Recipe1M to 1488. The employed query templates are provided in Figure 6.
Figure 7: Templates for Referring Segmentation task (One-to-Many).
Figure 8: Templates for Referring Segmentation task (One-to-Many/One-to-One/One-to-Zero).
Referring Segmentation To train our FoodLMM to generate masks for multiple objects in query texts and identify absent items in images, similar to [57], we establish data containing multi-target (One-to-Many), no-target (One-to-Zero), and single-target (One-to-One) expressions. We divide the query templates into two groups to accommodate these three scenarios: 1) We employ the templates listed in Figure 7 to create instruction-following data for One-to-Many scenario, where all templates guide the model to identify and segment all visible ingredients in food images. 2) Templates in Figure 8 are applicable to all three scenarios. For no-target data construction, $\{ingredient\_name\}$ represents one or multiple items absent in the image, and for multi-target and single-target data, it denotes one or several ingredients randomly selected from the image. $\{Is\_Are\}$ and $\{is\_are\}$ in the query text is automatically determined by the number of objects, to correctly use ’is’ or ’are’.
Figure 9: Templates for outlining the nutritional details of food and its ingredients.
Nutrition Estimation. To construct the nutrition estimation instruction-following data, we use the Nutrition5k dataset, which includes overhead RGB images and videos from various angles. Following the strategy described in [82], we sample every 5th frame from angle A and D videos for each dish. During training, for each dish, the probability of using the overhead RGB image is $70\%$ and the probability of randomly selecting a frame from the samples is $30\%$. We first establish three templates (Figure 9): 1) $ingredient\_calories$: the calorie content of ingredient (including weight and calories of the ingredient); 2) $ingredient\_nutrition$: the nutrition content of ingredient (including weight, calories, and three macro-nutrients); 3) $total\_nutrition$: the total nutrition content of the food (including the total weight, total calories, and the total content of the three macro-nutrients).
@@
Figure 11: Templates for querying calories of any ingredient in the food.
@@
Figure 13: Templates for querying about calories and the three macro-nutrient contents of any ingredient in the food.
Figure 15: Templates for querying the nutrient with the highest content among the three macro-nutrients in the food.
Figure 16: Templates for ranking the contents of the three macro-nutrients in the food.
We design seven question-and-answer patterns for nutrition estimation instruction-following data. These are as follows:
@@
2) Questions about calories of any ingredient in the food are presented in Figure 11, where $\{ingredient\_calories\}$ uses the template from Figure 9 and $\{ing\_names\}$ represents one or several randomly chosen ingredients;
3) Questions about the total calorie content and three macro-nutrients in the food, along with the calorie and macro-nutrient contents of each ingredient are outlined in Figure 12. The $\{ingredient\_nutrition\}$ and $\{total\_nutrition\}$ correspond to the templates provided in Figure 9;
4) Questions about calories and three macro-nutrient contents of any ingredient in the food are displayed in Figure 13. $\{ingredient\_nutrition\}$ refers to the template in Figure 9 and $\{ing\_names\}$ represents one or more randomly selected ingredients;
@@
6) Questions about the nutrient with the highest content among the three macro-nutrients in the food are shown in Figure 15, with $\{macronutrient\}$ indicating the nutrient with the highest content;
7) Ranking the contents of the three macro-nutrients in the food. Templates are displayed in Figure 16, where $\{macronutrient1\}$, $\{macronutrient2\}$, and $\{macronutrient3\}$ represent the nutrients with the highest, second-highest, and third-highest content respectively.
@@ 7.2 Stage 2: GPT-4 Generated Conversation Datasets
We send the detailed nutritional information of the food images (i.e., the total mass, calories, carbohydrates, fat, and protein of the food and these values of each ingredient) to GPT-4 and instruct GPT-4 to generate multi-round dialogues of specific topics. The prompt template used for calling the GPT-4 API is provided in Figure 17.
Figure 17: Prompt template for generating FoodDlialogues
In the prompt, we instruct GPT-4 to mark each nutritional element value with our Nutrition Estimation task-specific tokens and give an example to ensure that the dialogue generated by GPT-4 conforms to the format we need. We replace $\{DESC\}$, $\{TOPIC\}$ and $\{N\}$ in the template with the descriptions of different food images, the conversation topics that need to be generated, and the number of conversation turns respectively. For training FoodLMM, we take out the nutritional element values marked by task-specific tokens from the generated dialogue as supervision labels for the regression heads, while only task-specific tokens are retained in the text for training the LMM base. An example is shown in Figure 18.
@@
We send the ingredient lists of the selected images to GPT-4, requesting it to generate multi-round dialogues consisting of QA pairs that require complex reasoning. The prompt template is provided in Figure 19.
Figure 19: Prompt template for generating FoodReasonSeg
We ask GPT-4 to mark the ingredients mentioned in the answer with so that we can retrieve the mask labels from the original dataset. We also ask GPT-4 to generate segmentation tokens ([SEG]) in the answers for training FoodLMM. During training, we remove the braces while retaining the name of ingredient and segmentation token. An example is shown in Figure 20.
@@ 8.1 Implementation Details
Task-specific Tokens (Nutrition Estimation). We set 10 task-specific tokens corresponding to ten nutritional regression heads, We divide these 10 tokens into two categories: ingredient-level and dish-level. The ingredient-level tokens include: $\langle mass\rangle$, $\langle cal\rangle$, $\langle carb\rangle$, $\langle fat\rangle$, $\langle pro\rangle$ for regressing the mass, calories, carbohydrates, fats, and proteins of any ingredient respectively. And the dish-level tokens include: $\langle total\_mass\rangle$, $\langle total\_cal\rangle$, $\langle total\_carb\rangle$, $\langle total\_fat\rangle$, $\langle total\_pro\rangle$ for predicting the overall nutritional element values of the input food image. Note that in the implementation, we assigned numbers to the ingredient-level tokens, such as $\langle mass_{1}\rangle,\langle mass_{2}\rangle,\ldots,\langle mass_{n}\rangle$ to differentiate the nutritional elements of different ingredients mentioned in the answers. The LMM base (LLaVA) generates only these tokens without any specific value in response. Replace the corresponding tokens with the nutritional values obtained by regression heads to get the output answer.
Task-specific Tokens (Referring/Reasoning Segmentation). For referring and reasoning segmentation tasks, we set task-specific token $\langle seg\rangle$. The LMM base (LLaVA) learns to assign $\langle seg\rangle$ tokens to the ingredients mentioned in the answer, and maps its corresponding hidden state to hidden embedding through the MLP. Hidden embedding serves as the segmentation prompt. The decoder is trained to generate a segmentation mask according to a $\langle seg\rangle$ hidden embedding. Consistent with ingredient-level nutrition estimation tokens, $\langle seg\rangle$ tokens also use numerical numbers to distinguish different ingredients, i.e., $\langle seg_{1}\rangle,\langle seg_{2}\rangle,\ldots,\langle seg_{n}\rangle$. Summarize all generated masks to obtain the final segmented image containing multiple masks.
@@
Training/Fine-tuning details of different tasks. All of the fine-tuning experiments are conducted on 4 NVIDIA A100 (40G) GPUs. For Food Classification, the fine-tuning step on Food-101 is set to 2500, taking 13 hours on average with batch size 4. For Ingredient Recognition, the fine-tuning step on VIREO Food-172 is set to 2500, taking 9 hours with batch size 4. For Recipe Generation, the fine-tuning step on Recipe1M is set to 5000, taking 1 day with batch size 4. For Nutrition Estimation, the fine-tuning step on Nutrition5k is set to 2500, taking 10 hours with batch size 2. Different types of question-answer pairs in the instruction templates are selected randomly both in the training and fine-tuning. For referring nutrition estimation questions (in Figure 11 and Figure 13), we randomly select any number of ingredients from the image for the query, with the number of ingredients in each query capped at 20. For Referring Segmentation, FoodSeg103 and UECFoodPIXComplete are used together for fine-tuning. The step is set to 2500, taking 9 hours with batch size 2. The sampling rate of One-to-One/Many and One-to-Zero samples are set to 98% and 2% respectively both in the training and fine-tuning. For the FoodSeg103 dataset, the probabilities of sampling questions from templates in Figures 7 and Figure 8 are 0.4 and 0.6, respectively. For the UECFoodPixComplete dataset, since its annotations are not at ingredient-level and may not completely cover all ingredients, all the questions are sampled from templates in Figure 8. For one-to-many referring queries, we randomly select several ingredients from the image for the query, with the number of ingredients in each query capped at 20. For One-to-Zero referring queries, the maximum number of non-existent objects in a query is set to 3. The ratio of 1, 2 and 3 non-existent objects ($refer@1$, $refer@2$ and $refer@3$) in a One-to-Zero referring query is set to 2:1:1. To increase the diversity of the queries, non-existent referring is hybridized with existent referring, i.e., there may be both present ingredients and absent objects in the same query.
8.2 Additional results of the referring segmentation task.

### v2

@@ 5.3 FoodVQA Results
Our FoodVQA consists of three tasks, i.e., food classification, ingredient recognition, and recipe generation. For these tasks, we compare the performance of our model with baseline models that achieve state-of-the-art performances. Note that other LMMs (e.g., LISA [47]) are general-purpose with limited knowledge in the food domain, their outputs of the FoodVQA tasks are uncontrollable and difficult to be quantitatively evaluated. Empirically, the results are also not promising. Consequently, we do not include the results of other LMMs as baselines for the FoodVQA tasks. Table 3 summarizes the results. As shown in the table, for all these three tasks, our FoodLMM outperforms the SOTA methods and achieves the best results. Specifically, for food classification, by fine-tuning on Food-101 dataset, our model attains an accuracy of 93.93%, which is 3.07% higher than the SOTA method [69]. For ingredient recognition, our model outperforms CACLNet [59], a SOTA method on VIREO Food-172 dataset, for 3.2% in terms of F1. For recipe generation, without any extra information, our FoodLMM significantly outperforms FIRE [17], the SOTA method on Recipe1M dataset that leverages ground-truth ingredient labels for recipe generation. Figure 3) further shows qualitative examples for foodVQA. As can be seen, our FoodLMM can generate accurate answers for food classification, ingredient recognition as well as recipe generation, showing the effectiveness of our model.
@@ 5.4 Nutrition Estimation Results
We further evaluate the performance of our model in nutrition estimation, the results are shown in Table 3. After training Stage 1, FoodLMM has been able to accurately estimate the overall nutritional value based on food images, decreasing the prediction error by $4.5\%$ on average compared to the previous SOTA [83]. Examples of using FoodLMM to estimate the total nutritional value of food are shown in Figure 3. Moreover, FoodLMM is also able to predict the nutritional value of specific ingredients in dishes, which is named Referring Nutrition Estimation. It is worthwhile mentioning that previous methods lack the ability of Referring Nutrition Estimation. For evaluating the performance of Referring Nutrition Estimation, we select the top three ingredients with the largest mass as referred ingredients. Table 3 summarizes the results. In the table, refer@$k$st denotes referring the ingredient with the $k$th highest mass. The results show that although further fine-tuning does not enhance the performance of overall nutrition estimation, it significantly improves the accuracy of nutrition estimation for specific ingredients.
@@ 5.5 Referring Segmentation Results
We compare our FoodLMM with LISA [47] on FoodSeg103 and UECFoodPIX datasets, the results are shown in Table 3. For standard referring segmentation (one-to-one), FoodLMM’s cIoU scores significantly exceed those of LISA, achieving 20% improvement (0.65 to 0.78) on FoodSeg103 and 25% (0.72 to 0.90) on UECFoodPIX datasets. Since we deliberately train FoodLMM’s one-to-zero ability, the model may inevitably give a zero mask to the correct referring, i.e., mistakenly identifying the queried ingredient as non-existent. We examine the probability of accurate response (ACC) to measure this impact. Notably, the decrease in FoodLMM’s accuracy is minimal (99.16% on FoodSEG103 and 98.69% on UECFoodPIX) despite significant improvements in cIoU.
One-to-many. We then evaluate the performance of segmentation involving multiple ingredients referred in a query. Table  3 summarizes the segmentation results of three ($refer@3$) and five ($refer@5$) referred ingredients. The results show that FoodLMM, after the first training stage (FoodLMM S1), achieves high cIoU scores (around 0.7) on the $refer@3$ task on two datasets, while also maintaining high accuracy. However, as the number of referred ingredients increased the segmentation become more challenging. Nevertheless, the fine-tuned model (FoodLMM FT) attains higher cIoU scores across all scenarios and maintains a high probability of accurate responses (over 99%) even in complex segmentation tasks. Figure 3 shows the examples of referred ingredient segmentation.
One-to-zero To further examine the reasoning performance of our FoodLMM, we evaluate its performance under a challenging scenario, where the model is required to segment the ingredients that are not present in the image. For such a query, the model should reject to return the segmentation mask. The performances of one-to-zero segmentation are shown in Table 3. As LISA is not trained with one-to-zero samples, the accuracy in identifying absent ingredients on both FoodSeg103 and UECFoodPIX datasets is 0.0%. In contrast, our FoodLMM, attain an accuracy of 73.07% on FoodSeg103 and 77.50% on UECFoodPIX, when handling one single nonexistent referring ($refer@1$), after the first train stage. The task becomes more challenging when there are three absent referring (refer@3) in the query. Nevertheless, FoodLMM still maintains a high accuracy, surpassing 80% after fine-tuning. Figure 3 presents an example where the user requests FoodLMM to segment watermelon from the seafood vegetable soup, but the model declines.
@@ 5.6 Performance on the proposed Benchmarks
Multi-turn Conversation. We further evaluate the performance of our model on the Multi-turn Conversation dataset. The evaluation is performed by user study. Specifically, we invited 10 participants to manually evaluate the satisfaction score (1 to 5) of 20 sets of multi-turn dialogues (totalling 200 dialogue sets, all of the questions are randomly selected from the test set of our FoodDialogues). The results are summarized in Table 4. The majority of the dialogues (173 out of 200) receive high scores (>>>=3), and 4 is the most frequent score. It indicates that our FoodLMM is capable of closely aligning with user preferences and generating high-quality answers that are both accurate and explanatory.
@@
Reasoning Segmentation. We compare the performance of our FoodLMM with LISA in reasoning segmentation on our FoodReasonSeg benchmark. The results are shown in Table 5. From the results, LISA achieves a quite low gIoU and cIoU, which are $0.19$ and $0.17$, respectively. It indicates that although LISA is excellent in the general domain’s reasoning segmentation, due to lacking specialized knowledge in the food domain, its performance in food reasoning segmentation is less impressive. In contrast, our FoodLMM S1 achieves much better performance by learning food knowledge from public datasets. After the Stage 2 fine-tuning, the performance of FoodLMM is significantly improved, with the gIoU score increasing from $0.35$ to $0.71$ and the cIoU score from $0.34$ to $0.72$ compared to FoodLMM S1, indicates that FoodLMM acquired more expertise and stronger reasoning ability from the constructed datasets.
5.7 Discussion about Multi-task and Single-task learning
Referring Segmentation One-to-One/Many (cIoU/Acc) Referring Segmentation One-to-Zero (Acc)
Multitask 0.75 / 99.25% Multitask 86.02%
Single-task 0.76 / 98.20% Single-task 89.07%
Nutrition Estimation (MAE) Referring Nutrition Estimation (MAE)
Multitask 24.48 / 28.8% Multitask 10.64 / 34.60%
Single-task 25.6 / 30.5% Single-task 16.40 / 43.62%
Ingredient Recognition (IoU/F1) Recipe Generation (BLEU/Rouge-L)
Multitask 56.94 / 68.97 Multitask 5.89 / 35.12
Single-task 56.49 / 68.72 Single-task 6.24 / 36.65
Table 6: Performance comparison between multi-task learning and single-task learning. The average results are reported.
We compare the performance between single-task and multi-task learning of our FoodLMM in this section. Multi-task model is capable of dealing with various food tasks, while single-task model focuses on a specific task. Table 6 lists the performance differences between multi-task and single-task models. For the Referring Segmentation task, we present average segmentation results for variable number of ingredients across FoodSeg103 [96] and UEC-FoodPIX Complete [73]. For Nutrition Estimation task, we report the average prediction results for five essential nutrients. We observe that the advantage of multi-task is the ability to leverage knowledge across different tasks, which benefits some basic recognition tasks such as ingredient recognition and nutrition estimation. For more complex tasks, multi-task is only able to boost the performance of referring nutrition estimation. In general, single-task learning may benefit complex tasks such as recipe generation as model training can be more easily optimized. Nevertheless, multi-task learning still suffers from the problem of gradient conflicts between tasks [58, 98, 19], which remains an open issue and is worth further research.
@@ 7.1 Stage 1: Public Food Datasets
Figure 4: Templates for Food Classification and Ingredient Recognition tasks.
Food Classification. We adopt the VIREO Food-172 to construct the instruction-following data for the food classification task. Our 10 question templates are detailed on the left of Figure 4.
Ingredient Recognition. We use VIREO Food-172 and Recipe1M to develop instruction-following data for the ingredient recognition task. The templates are outlined on the right of Figure 4.
Recipe Generation. We construct instruction-following data for the recipe generation task using recipes from the Recipe1M dataset that have corresponding images. Moreover, consistent with [80], we map the 16k ingredients in Recipe1M to 1488. The employed query templates are provided on the left of Figure 5.
Figure 5: Templates for Recipe Generation task and outlining the nutritional details of food and its ingredients.
Figure 6: Templates for querying calories of any ingredient in the food and querying about calories and the three macro-nutrient contents of any ingredient in the food.
Nutrition Estimation. To construct the nutrition estimation instruction-following data, we use the Nutrition5k dataset, which includes overhead RGB images and videos from various angles. Following the strategy described in [83], we sample every 5th frame from angle A and D videos for each dish. During training, for each dish, the probability of using the overhead RGB image is $70\%$ and the probability of randomly selecting a frame from the samples is $30\%$. We first establish three templates (on the right of Figure 5): 1) $ingredient\_calories$: the calorie content of ingredient (including weight and calories of the ingredient); 2) $ingredient\_nutrition$: the nutrition content of ingredient (including weight, calories, and three macro-nutrients); 3) $total\_nutrition$: the total nutrition content of the food (including the total weight, total calories, and the total content of the three macro-nutrients).
@@
Figure 10: Templates for querying the nutrient with the highest content among the three macro-nutrients and ranking the contents of the three macro-nutrients in the food.
We design seven question-answer patterns for nutrition estimation instruction-following data. These are as follows:
@@
2) Questions about calories of any ingredient in the food are presented on the left of Figure 6, where $\{ingredient\_calories\}$ uses the template from Figure 5 and $\{ing\_names\}$ represents one or several randomly chosen ingredients;
3) Questions about the total calorie content and three macro-nutrients in the food, along with each ingredient’s calorie and macro-nutrient contents are outlined in Figure 9. The $\{ingredient\_nutrition\}$ and $\{total\_nutrition\}$ correspond to the templates provided in Figure 5;
4) Questions about calories and three macro-nutrient contents of any ingredient in the food are displayed on the right of Figure 6. $\{ingredient\_nutrition\}$ refers to the template in Figure 5 and $\{ing\_names\}$ represents one or more randomly selected ingredients;
@@
6) Questions about the nutrient with the highest content among the three macro-nutrients in the food are shown on the left of Figure 10, with $\{macronutrient\}$ indicating the nutrient with the highest content;
7) Ranking the contents of the three macro-nutrients in the food. Templates are displayed on the right of Figure 10, where $\{macronutrient1\}$, $\{macronutrient2\}$, and $\{macronutrient3\}$ represent the nutrients with the highest, second-highest, and third-highest content respectively.
Figure 11: Templates for Referring Segmentation task.
Referring Segmentation. To train our FoodLMM to generate masks for multiple objects in query texts and identify absent items in images, similar to [55], we establish data containing multi-target (One-to-Many), no-target (One-to-Zero), and single-target (One-to-One) expressions. We divide the query templates into two groups to accommodate these three scenarios: 1) We employ the templates listed on the left of Figure 11 to create instruction-following data for the One-to-Many scenario, where all templates guide the model to identify and segment all visible ingredients in food images. 2) Templates on the right of Figure 11 are applicable to all three scenarios. For no-target data construction, $\{ingredient\_name\}$ represents one or multiple items absent in the image, and for multi-target and single-target data, it denotes one or several ingredients randomly selected from the image. $\{Is\_Are\}$ and $\{is\_are\}$ in the query text is automatically determined by the number of objects, to correctly use ’is’ or ’are’.
@@ 7.2 Stage 2: GPT-4 Generated Conversation Datasets
Figure 12: Prompt template for generating FoodDlialogues.
We send the detailed nutritional information of the food images (i.e., the total mass, calories, carbohydrates, fat, and protein of the food and these values of each ingredient) to GPT-4 and instruct GPT-4 to generate multi-round dialogues of specific topics. The prompt template used for calling the GPT-4 API is provided in Figure 12. In the prompt, we instruct GPT-4 to mark each nutritional element value with our Nutrition Estimation task-specific tokens and give an example to ensure that the dialogue generated by GPT-4 conforms to the format we need. We replace $\{DESC\}$, $\{TOPIC\}$ and $\{N\}$ in the template with the descriptions of different food images, the conversation topics that need to be generated, and the number of conversation turns respectively. For training FoodLMM, we take out the nutritional element values marked by task-specific tokens from the generated dialogue as supervision labels for the regression heads, while only task-specific tokens are retained in the text for training the LMM base. An example is shown in Figure 13.
@@
We send the ingredient lists of the selected images to GPT-4, requesting it to generate multi-round dialogues consisting of QA pairs that require complex reasoning. The prompt template is provided in Figure 14. We ask GPT-4 to mark the ingredients mentioned in the answer with so that we can retrieve the mask labels from the original dataset. We also ask GPT-4 to generate segmentation tokens ([SEG]) in the answers for training FoodLMM. During training, we remove the braces while retaining the name of ingredient and segmentation token. An example is shown in Figure 15.
Figure 14: Prompt template for generating FoodReasonSeg.
@@ 8.1 Implementation Details
[…]
@@
Training/Fine-tuning details of different tasks. All of the fine-tuning experiments are conducted on 4 NVIDIA A100 (40G) GPUs. For Food Classification, the fine-tuning step on Food-101 is set to 2500, taking 13 hours on average with batch size 4. For Ingredient Recognition, the fine-tuning step on VIREO Food-172 is set to 2500, taking 9 hours with batch size 4. For Recipe Generation, the fine-tuning step on Recipe1M is set to 5000, taking 1 day with batch size 4. For Nutrition Estimation, the fine-tuning step on Nutrition5k is set to 2500, taking 10 hours with batch size 2. Different types of question-answer pairs in the instruction templates are selected randomly both in the training and fine-tuning. For referring nutrition estimation questions (in Figure 6), we randomly select any number of ingredients from the image for the query, with the number of ingredients in each query capped at 20. For Referring Segmentation, FoodSeg103 and UECFoodPIXComplete are used together for fine-tuning. The step is set to 2500, taking 9 hours with batch size 2. The sampling rate of One-to-One/Many and One-to-Zero samples are set to 98% and 2% respectively both in the training and fine-tuning. For the FoodSeg103 dataset, the probabilities of sampling questions from templates in Figures 11 and Figure 11 are 0.4 and 0.6, respectively. For the UECFoodPixComplete dataset, since its annotations are not at ingredient-level and may not completely cover all ingredients, all the questions are sampled from templates in Figure 11. For one-to-many referring queries, we randomly select several ingredients from the image for the query, with the number of ingredients in each query capped at 20. For One-to-Zero referring queries, the maximum number of non-existent objects in a query is set to 3. The ratio of 1, 2 and 3 non-existent objects ($refer@1$, $refer@2$ and $refer@3$) in a One-to-Zero referring query is set to 2:1:1. To increase the diversity of the queries, non-existent referring is hybridized with existent referring, i.e., there may be both present ingredients and absent objects in the same query.
8.2 Additional results of the referring segmentation task